Circular No. 7391

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE Further to IAUC 7391, B. E. Schaefer reports eleven additional appar-

ent supernovae. SN 2000ba has no apparent host galaxy, to R ∼ 21.5.

SN  2000 UT  α_{2000}  δ_{2000}  R  Offset
2000av Mar. 10  7^{h}38^{m}28.^{s}85 − 1^{o}44′16′′.0 20.1 0′′ E, 1′′ S
2000aw Mar. 10 8 01 34.00 − 1 57 14.3 19.4 0′′ E, 0′′ N
2000ax Mar. 10 8 44 16.35 − 0 43 02.7 20.6 0′′ E, 1′′ S
2000ay Mar. 10 9 05 44.86 − 1 39 52.8 19.8 0′′ E, 1′′ N
2000az Mar. 10 9 18 02.15 − 2 16 58.9 19.8 0′′ E, 1′′ N
2000ba Mar. 9 10 16 02.05 − 1 53 39.9 19.9  —
2000bb Mar. 1 10 32 38.83 − 2 00 24.2 20.2 1′′ E, 0′′ N
2000bc Mar. 9 10 33 06.30 − 1 00 21.2 18.9 3′′ E, 2′′ S
2000bd Mar. 10 13 07 20.81 − 1 05 08.1 20.5 1′′ W, 1′′ S
2000be Mar. 9 15 12 03.38 − 0 00 04.1 20.1 0′′ E, 0′′ N
2000bf Mar. 9 15 12 18.37 − 1 39 17.5 20.7 0′′ E, 0′′ N

SUPERNOVA 1998bu IN NGC 3368 E. Cappellaro, S. Benetti, A. Pastorello, M. Turatto, G. Altavilla, and

L. Rizzi, Padova Observatory, write: “We observed SN 1998bu (IAUC 6899, 6905; Jha et al. 1999, Ap.J. Suppl. 125, 73, and references therein) on Feb. 15 with the Danish 1.5-m telescope (+ DFOSC) and on Mar. 2 with the European Southern Observatory 3.6-m telescope (+ EFOSC2) at La Silla. For the first date, corresponding to 637 days after maximum, we measured B = 21.07 and V = 21.16. This indicates that the decline had slowed, the new points being ∼ 3 mag brighter than the extrapolated exponential tail. The spectrum taken on Mar. 2 (day 653) show a blue energy distribution with broad absorption and emission features. The highest intensity peaks are centered at 355, 394, 408, and 460 nm. Both light curve and spectrum show an impressive similarity to SN 1991T at the same phase. Since the late pecularities of SN 1991T were convincingly interpreted as due to reprocessing of the luminosity at maximum by foreground dust (Schmidt et al. 1994, Ap.J. 434, L19), the same interpretation may apply also to this case. It may not be a coincidence that the early luminosity decline rates of the two supernovae were also very similar.”

2000 March 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green